OpenStudy (anonymous):

suppose that a particle moves along the x-axis so that at time t, the particle is located xt=2t^3-4t^2 meters to the right of the origin. Find the average velocity of the particle between t=4 and t=5. Find the average velocity of the particle between t=4 and t=4.01. Find the instantaneous velocity of the particle at t=4.

OpenStudy (anonymous):

The average value of a function A is given by: \[\frac{1}{b-a}\int\limits_a^b A dt \] The velocity is given by: \[\frac{d}{dt}(x(t))=v(t)\] if x(t) is the position. The instantaneous velocity at a time t_0 is \[\frac{d}{dt}x(t) |_{t=t_0}=v(t_0) \]

OpenStudy (agreene):

they give you the displacement vector: r=2t^3-4t^2 from this you need to find velocities: so remember r'=v(t) r'=d/dt (2t^3-4t^2) r'=v(t)=6t^2-8t now that we have velocity, plug in for the needed times and average where needed and such.